OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 2y = 6 x + 3y = −1 −2 2 5 −5

OpenStudy (anonymous):

thought this one was b but i don't think it goes to the second equation

OpenStudy (anonymous):

6x+6y=18 -2x-6y=2 4x = 20

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i multiply the top by 3 and the bottom by -2 so i can eliminate the y leaves you with only the x

OpenStudy (anonymous):

multiply the first equation by 3 and the second equation by 2 the first equation becomes : 6x+6y =18 the second equation becomes : 2x+6y= -2 on subtracting first equation from second : 4x=20 => x=20/4 =5 => x=5

OpenStudy (anonymous):

ohh I see what you are both are saying

OpenStudy (anonymous):

 this is what I got